# The Distributional Uncertainty of the SHAP score in Explainable Machine Learning

### v1

@@ Abstract
Attribution scores reflect how important the feature values in an input entity are for the output of a machine learning model. One of the most popular attribution scores is the SHAP score, which is an instantiation of the general Shapley value used in coalition game theory. The definition of this score relies on a probability distribution on the entity population. Since the exact distribution is generally unknown, it needs to be assigned subjectively or be estimated from data, which may lead to misleading feature scores. In this paper, we propose a principled framework for reasoning on SHAP scores under unknown entity population distributions. In our framework, we consider an uncertainty region that contains the potential distributions, and the SHAP score of a feature becomes a function defined over this region. We study the basic problems of finding maxima and minima of this function, which allows us to determine tight ranges for the SHAP scores of all features. In particular, we pinpoint the complexity of these problems, and other related ones, showing them to be NP-complete. Finally, we present experiments on a real-world dataset, showing that our framework may contribute to a more robust feature scoring.
@@ 1 Introduction
Proposing and investigating different forms of explaining and interpreting the outcomes from AI-based systems has become an effervescent area of research and applications Molnar (2020), leading to the emergence of the area of Explainable Machine Learning. In particular, one wants to explain results obtained from ML-based classification systems. A widespread approach to achieve this consists in assigning numerical attribution scores to the feature values that, together, represent a given entity under classification, for which a given label has been obtained. The score of a feature value indicates how relevant it is for this output label.
One of the most popular attribution scores is the so-called SHAP score Lundberg and Lee (2017); Lundberg, S. et al. (2020), which is a particular form of the general Shapley value used in coalition game theory Shapley (1953); Roth (1988). SHAP, as every instantiation of the Shapley value, requires a wealth function shared by all the players in the coalition game. SHAP uses one that is an expected value based on a probability distribution on the entity population. [^1] Since SHAP’s inception, several variations have been proposed and investigated, but they all rely on some probability distribution. In this work we stick to the original formulation. Since the exact distribution is generally unknown, it needs to be assigned subjectively or be estimated from data. This may lead to different kinds of errors, and in particular, to misleading feature scores.
In this work, we propose a principled framework for reasoning on SHAP scores under distributional uncertainty, that is, under an unknown distribution over the entity population. We focus on binary classifiers, i.e., classifiers that returns $1$ (accept) or 00 (reject). We also assume that the inputs to these classifiers are binary features, i.e., features that can take values 00 or $1$. Furthermore, we focus on product distributions. Their use for SHAP computation is common, and imposes feature independence Arenas et al. (2023); Van Den Broeck et al. (2022). In practice, one frequently uses an empirical product distribution (of the empirical marginals), which may vary depending on the data set from which the sampling is performed Bertossi et al. (2020). We see our concentration on product distributions as an important first step towards the distributional analysis of SHAP.
@@
Consider the classifier $M$ given in Table 1. Let $e$ be the null entity (first row), and assume a product distribution $\langle p_{x},p_{y},p_{z}\rangle$ over the feature space, e.g. $\mathbb{P}(x=1,y=0,z=1)=p_{x}(1-p_{y})p_{z}$. The SHAP score for entity $e$ and feature $z$ depends on the probabilities $p_{x}$, $p_{y}$ and $p_{z}$, and this relation can be expressed through the following function:
$Shap(M,e,z)=Shap_{M,e,z}(p_{x},p_{y},p_{z})$
$=\frac{1}{6}p_{z}(-4p_{x}p_{y}+3p_{x}+3p_{y}),$ (1)
@@
To determine the SHAP intervals it is necessary to find minimal upper-bounds and maximal lower-bounds for the SHAP score in the uncertainty region. Formulated in terms of thresholds, these problems turn out to be in the class NP for a wide class of classifiers. Furthermore, we establish that this problem becomes NP-complete even for simple models such as decision trees (and other classifiers that share some properties with them). Notice that computing SHAP for decision trees can be done in polynomial time under the product and uniform distributions. Actually, this result can be obtained for larger classes of classifiers that include decision trees Arenas et al. (2023); Van Den Broeck et al. (2022).
We also propose and study three other problems related to the behavior of the SHAP score in the uncertainty region, and obtain the same complexity theoretical results as for the problem of computing the maximum and minimum SHAP score. These problems are: (1) deciding whether there are two different distributions in the uncertainty region such that the SHAP score is positive in one of them and negative in the other one (and therefore there is no certainty on whether the feature contributes positively or negatively to the prediction), (2) deciding if there is some distribution such that the SHAP score is 0 (i.e. if the feature can be considered irrelevant in some sense), and (3) deciding if for every distribution in the uncertainty region it holds that a feature $x$ is better ranked than a feature $y$ (i.e. if $x$ dominates $y$).
@@
Close to our work, but aiming towards a different direction, we find the problem of distributional shifts Carvalho et al. (2019); Molnar (2020); Linardatos et al. (2020), which in ML occur when the distribution of the training data differs from the data the classifier encounters in a particular application. This discrepancy poses significant challenges, as can lead to decreased performance and unexpected behavior of models.
Also related is the problem of score robustness Alvarez-Melis and Jaakkola (2018); Huang and Marques-Silva (2023): one can analyze how scores change under small perturbations of an input. In our case, we study uncertainty at the level of the underlying probability distributions.
The work Slack et al. (2021) also tries to address uncertainty in the importance of features for local explanations, but does so from a Bayesian perspective: they use a novel sampling procedure to estimate credible intervals around the mean of the feature importance, and derive closed-form expressions for the number of perturbations required to reach explanations within the desired levels of confidence.
Finding optimal intervals under uncertainty as we do here, is reminiscent of finding tight ranges for aggregate queries from uncertain databases which are repaired due to violations of integrity constraints Arenas et al. (2003).
@@
This paper is structured as follows:  In Section 2 we introduce notation and recall basic definitions. In Section 3, we formalize our problems and obtain the first results. Section 4 presents our main complexity results. In Section 5, we describe our experiments and show their outcome. Finally, in Section 6, we make some final remarks and point to open problems. Full proofs can be found in the Appendix.
@@ 2 Preliminaries
A (binary) classifier or model $M$ over $X$ is a mapping $M:\texttt{ent}(X)\to\{0,1\}$. We say that $M$ accepts $e$ if $M(e)=1$, otherwise $M$ rejects the entity. Let $M$ be a Boolean classifier and $e$ an entity, both over $X$. We define the function $\phi_{M,e}:2^{X}\to[0,1]$ as:
@@
The function $\phi_{M,e}$ can be used as the wealth function in the general formula of the Shapley value Shapley (1953); Roth (1988) to obtain the SHAP score of the feature values in $e$.
@@
Intuitively, the SHAP score intends to measure how the inclusion of $x$ affects the conditional expectation of the prediction. In order to do this, it considers every possible subset $S\subseteq X\setminus\{x\}$ of the features and compares the expectation for the set $S$ against $S\cup\{x\}$. A score close to $1$ implies that $x$ heavily leans the classifier $M$ towards acceptance, while a score close to $-1$ indicates that it leans the prediction towards rejection (note SHAP always takes values in $[-1,1]$).
@@
In practical applications, the exact distribution on entities is generally unknown and subjectively assumed or estimated from data. The previous example shows that the choice of the underlying distribution can have severe effects when establishing the importance of features in the classifications. To overcome these problems, in the next section we formalize the notion of distributional uncertainty and present our framework for reasoning about SHAP scores in that setting.
@@ 3 SHAP under Distributional Uncertainty
The general idea of our framework is as follows: we explicitly consider a set that contains the potential distributions. This provide us with what we call the uncertainty region. This allows us to reinterpret the SHAP score as a function from the uncertainty region to $\mathbb{R}$. We can then analyze the behavior of this function in order to gain concrete insights about the importance of a feature.
Recall a product distribution is determined by its parameters $(p_{x})_{x\in X}$. For convenience, we always assume an implicit ordering on the features. Hence, we can identify our space of probability distributions over $\texttt{ent}(X)$ with the set $[0,1]^{|X|}$. In order to define uncertainty regions, it is natural then to consider hyperrectangles $\mathcal{I}\subseteq[0,1]^{|X|}$, i.e., subsets of the form =[$\mathcal{I}={}_{x\in X}[a_{x},b_{x}]$. Intuitively, these regions correspond to independently choosing a confidence interval $[a_{x},b_{x}]$ for the unknown probability $p_{x}$, for each feature $x\in X$.
@@
We can then reason about the importance of features via the analysis of SHAP polynomials. Here we concentrate on the fundamental problems of finding maxima and minima of these polynomials over the uncertainty region. This allows us to determine the SHAP interval of a feature, i.e. the set of possible SHAP scores of the feature over the uncertainty region. SHAP intervals may provide useful insights. For instance, obtaining smaller SHAP intervals for a feature suggests its SHAP score is more robust against uncertainty. On the other hand, they can be used to assess the relative importance of features (see Section 5 for more details).
@@
Problem: REGION-MAX-SHAP Input: A classifier $M$, an entity $e$, a feature $x$, a hyperrectangle $\mathcal{I}$ and a rational number $b$. Output: Is there a point $\textbf{p}\in\mathcal{I}$ such that $Shap_{M,e,x}(\textbf{p})\geq b$?.
The problem REGION-MIN-SHAP is defined analogously by requiring $Shap_{M,e,x}(\textbf{p})\leq b$. We also consider some related problems:
@@ 4.1 Preliminaries on multilinear polynomials
A vertex of a hyperrectangle =×$\mathcal{I}=\times_{i=1}^{n}[a_{i},b_{i}]$ is a point $\textbf{p}\in\mathcal{I}$ such that $\textbf{p}_{i}\in\{a_{i},b_{i}\}$ for each $1\leq i\leq n$. The following is a well-known fact about multilinear polynomials (see e.g. Laneve et al. (2010)). For completeness we provide a simple self-contained proof in Appendix A.1.
@@
Proposition 4 yields two algorithmic consequences. On the one hand, it induces an algorithm to find the maximum of $f$ over $\mathcal{I}$ in time $2^{n}poly(|f|)$: simply evaluate the polynomial on all the vertices, and keep the maximum [^3] We assume $f$ is given by listing its non-zero coefficients and their corresponding monomials.. On the other hand, it shows that this problem is certifiable: to decide whether $f$ can reach a value as big as $b$, we just need to guess the corresponding vertex and evaluate it. We show that, within the usual complexity theoretical assumptions, there is no polynomial algorithm to find this maximum (see Appendix A.1):
@@
Given a multilinear polynomial $f$, a hyperrectangle =[$\mathcal{I}={}_{i=1}^{n}[a_{i},b_{i}]$, and a rational $b$, the problem of deciding whether there is an $\textbf{x}\in\mathcal{I}$ such that $f(\textbf{x})\geq b$ is NP-complete.
@@ 4.2 Complexity of REGION-MAX-SHAP
It is well-known that computing SHAP scores is hard for general classifiers. For instance, the problem is already #P-hard when considering Boolean circuits Arenas et al. (2023) or logistic regression models Van Den Broeck et al. (2022). For linear perceptrons, model counting is intractable Barceló et al. (2020) and by the results in Arenas et al. (2023), it follows that SHAP computation for perceptrons is also intractable.  On the other hand, computing maxima of SHAP polynomials for a certain class of classifiers is as hard as computing SHAP scores for that class of classifiers: if we consider the hyperrectangle consisting of a single point =[$\mathcal{I}={}_{i=1}^{n}[p_{i},p_{i}]$, the maximum of the SHAP polynomial coincides with the SHAP score for the product distribution $(p_{i})_{1\leq i\leq n}$. Therefore, we focus on family of classifiers where the SHAP score can be computed in polynomial time. A prominent example is the class of decomposable and deterministic Boolean circuits, whose tractable SHAP score computation has been shown recently Arenas et al. (2023). This class contains as a special case the well-known class of decision trees. For a formal definition see  Darwiche (2001); Darwiche and Marquis (2002). As a consequence of Proposition 4, we obtain the following complexity upper bound:
@@
Membership in NP follows from Corollary 6. For the hardness part, we reduce from the well-known NP-complete problem VERTEX-COVER: given a graph $G=(V,E)$ and $k\geq 1$, decide whether there is a vertex cover [^4] Recall a vertex cover of $G=(V,E)$ is a subset of the nodes $C\subseteq V$ such that for each edge $\{u,v\}\in E$, either $u\in C$ or $v\in C$. of size at most $k$.
@@
where $\ell$ is the size of the subset $C$ and the term $T_{n,\ell}$ grows as $\ell$ grows. The term $I_{u,v}$ is positive and works as a penalization factor which is “activated” if $p_{u}=p_{v}=1$, i.e. when the edge $uv$ is uncovered. Furthermore, by adding an extra feature $w$ to the model (and consequently, another probability $p_{w}$) we can make this penalization factor arbitrarily big in relation to the size factor $T_{n,\ell}$.
We choose the bound $b$ to be $-T_{n,k}$. If $C$ is a vertex cover of size $\ell$, then each term $p_{u}p_{v}I_{u,v}$ equals 0 and hence $Shap_{M,e,x}(\textbf{p}^{C})=-T_{n,\ell}$. On the other hand, if $C$ is not a vertex cover, then some $p_{u}p_{v}I_{u,v}$ is non-zero, and by defining an adequate interval for $p_{w}$ we can ensure that $-p_{v}p_{v}I_{u,v}<b$. Hence the only way to obtain $Shap_{M,e_{0},x_{0}}(\textbf{p}^{C})\geq b$ is to pick $C$ to be, in the first place, a vertex cover, and secondly, one of size $\ell\leq k$. ∎

### v2

@@ Abstract
Attribution scores reflect how important the feature values in an input entity are for the output of a machine learning model. One of the most popular attribution scores is the SHAP score, which is an instantiation of the general Shapley value used in coalition game theory. The definition of this score relies on a probability distribution on the entity population. Since the exact distribution is generally unknown, it needs to be assigned subjectively or be estimated from data, which may lead to misleading feature scores. In this paper, we propose a principled framework for reasoning on SHAP scores under unknown entity population distributions. In our framework, we consider an uncertainty region that contains the potential distributions, and the SHAP score of a feature becomes a function defined over this region. We study the basic problems of finding maxima and minima of this function, which allows us to determine tight ranges for the SHAP scores of all features. In particular, we pinpoint the complexity of these problems, and other related ones, showing them to be intractable. Finally, we present experiments on a real-world dataset, showing that our framework may contribute to a more robust feature scoring.
[…]
123
@@ 1 Introduction
Proposing and investigating different forms of explaining and interpreting the outcomes from AI-based systems has become an effervescent area of research and applications [18], leading to the emergence of the area of Explainable Machine Learning. In particular, one wants to explain results obtained from ML-based classification systems. A widespread approach to achieve this consists in assigning numerical attribution scores to the feature values that, together, represent a given entity under classification, for which a given label has been obtained. The score of a feature value indicates how relevant it is for this output label.
One of the most popular attribution scores is the so-called SHAP score [16, 22], which is a particular form of the general Shapley value used in coalition game theory [23, 21]. SHAP, as every instantiation of the Shapley value, requires a wealth function shared by all the players in the coalition game. SHAP uses one that is an expected value based on a probability distribution on the entity population. [^2] Since SHAP’s inception, several variations have been proposed and investigated, but they all rely on some probability distribution. In this work we stick to the original formulation. Since the exact distribution is generally unknown, it needs to be assigned subjectively or be estimated from data. This may lead to different kinds of errors, and in particular, to misleading feature scores.
In this work, we propose a principled framework for reasoning on SHAP scores under distributional uncertainty, that is, under an unknown distribution over the entity population. We focus on binary classifiers, i.e., classifiers that returns $1$ (accept) or 00 (reject). We also assume that the inputs to these classifiers are binary features, i.e., features that can take values 00 or $1$. Furthermore, we focus on product distributions. Their use for SHAP computation is common, and imposes feature independence [3, 27]. In practice, one frequently uses an empirical product distribution (of the empirical marginals), which may vary depending on the data set from which the sampling is performed [7]. We see our concentration on product distributions as an important first step towards the distributional analysis of SHAP.
@@
Consider the classifier $M$ given in Table 1. Let $e$ be the null entity (first row), and assume a product distribution $\langle p_{x},p_{y},p_{z}\rangle$ over the feature space, e.g., $\mathbb{P}(x=1,y=0,z=1)=p_{x}(1-p_{y})p_{z}$. The SHAP score for entity $e$ and feature $z$ depends on the probabilities $p_{x}$, $p_{y}$ and $p_{z}$, and this relation can be expressed through the following function:
$Shap(M,e,z)$ $=$ $Shap_{M,e,z}(p_{x},p_{y},p_{z})$ (1)
$=$ $\frac{1}{6}p_{z}(-4p_{x}p_{y}+3p_{x}+3p_{y}).$
@@
To determine the SHAP intervals it is necessary to find minimal upper-bounds and maximal lower-bounds for the SHAP score in the uncertainty region. Formulated in terms of thresholds, these problems turn out to be in the class NP [^3] See [12] for a standard introduction into the complexity classes considered in this paper. for a wide class of classifiers. Furthermore, we establish that this problem becomes NP-complete even for simple models such as decision trees (and other classifiers that share some properties with them). Notice that computing SHAP for decision trees can be done in polynomial time under the product and uniform distributions. Actually, this result can be obtained for larger classes of classifiers that include decision trees [3, 27].
We also propose and study three other problems related to the behavior of the SHAP score in the uncertainty region, and obtain the same complexity theoretical results as for the problem of computing the maximum and minimum SHAP score. These problems are: (1) deciding whether there are two different distributions in the uncertainty region such that the SHAP score is positive in one of them and negative in the other one (and therefore there is no certainty on whether the feature contributes positively or negatively to the prediction), (2) deciding if there is some distribution such that the SHAP score is 0 (i.e., if the feature can be considered irrelevant in some sense), and (3) deciding if for every distribution in the uncertainty region it holds that a feature $x$ is better ranked than a feature $y$ (i.e., if $x$ dominates $y$).
@@
Close to our work, but aiming towards a different direction, we find the problem of distributional shifts [8, 18, 15], which in ML occur when the distribution of the training data differs from the data the classifier encounters in a particular application. This discrepancy poses significant challenges, as can lead to decreased performance and unexpected behavior of models.
Also related is the problem of score robustness [1, 13]: one can analyze how scores change under small perturbations of an input. In our case, we study uncertainty at the level of the underlying probability distributions.
The work [25] also tries to address uncertainty in the importance of features for local explanations, but does so from a Bayesian perspective: they use a novel sampling procedure to estimate credible intervals around the mean of the feature importance, and derive closed-form expressions for the number of perturbations required to reach explanations within the desired levels of confidence.
Finding optimal intervals under uncertainty as done here, is reminiscent of finding tight ranges for aggregate queries from uncertain databases which are repaired due to violations of integrity constraints [2].
Finally, other lines of work such as [17] aim to understand the uncertainty that arises from approximation errors when computing the Shapley values via a sampling procedure over the feature space. In such contexts, the distribution is usually assumed as given, and therefore these works focus on formalizing a scenario that differs from ours. Moreover, all our analyses and algorithms are based on optimal confidence intervals that arise from exact computation of the Shapley values.
@@
This paper is structured as follows:  In Section 2 we introduce notation and recall basic definitions. In Section 3, we formalize our problems and obtain the first results. Section 4 presents our main complexity results. In Section 5, we describe our experiments and show their outcome. Finally, in Section 6, we make some final remarks and point to open problems. Full proofs can be found in the arXiv version of the paper [9].
@@ 2 Preliminaries
A (binary) classifier or model $M$ over $X$ is a mapping $M:\texttt{ent}(X)\to\{0,1\}$. We say that $M$ accepts $e$ if $M(e)=1$, otherwise $M$ rejects the entity. Let $M$ be a binary classifier and $e$ an entity, both over $X$. We define the function $\phi_{M,e}:2^{X}\to[0,1]$ as:
@@
The function $\phi_{M,e}$ can be used as the wealth function in the general formula of the Shapley value [23, 21] to obtain the SHAP score of the feature values in $e$.
@@
Intuitively, the SHAP score intends to measure how the inclusion of $x$ affects the conditional expectation of the prediction. In order to do this, it considers every possible subset $S\subseteq X\setminus\{x\}$ of the features and compares the expectation for the set $S$ against $S\cup\{x\}$. A score close to $1$ implies that $x$ heavily leans the classifier $M$ towards acceptance, while a score close to $-1$ indicates that it leans the prediction towards rejection (note that SHAP always takes values in $[-1,1]$).
@@
In practical applications, the exact distribution of entities is generally unknown and subjectively assumed or estimated from data. The previous example shows that the choice of the underlying distribution can have severe effects when establishing the importance of features in the classifications. To overcome these problems, in the next section we formalize the notion of distributional uncertainty and present our framework for reasoning about SHAP scores in that setting.
@@ 3 SHAP under Distributional Uncertainty
The general idea of our framework is as follows: we explicitly consider a set that contains the potential distributions. This provides us with what we call the uncertainty region. This allows us to reinterpret the SHAP score as a function from the uncertainty region to $\mathbb{R}$. We can then analyze the behavior of this function in order to gain concrete insights about the importance of a feature.
Recall that a product distribution is determined by its parameters $(p_{x})_{x\in X}$. For convenience, we always assume an implicit ordering on the features. Hence, we can identify our space of probability distributions over $\texttt{ent}(X)$ with the set $[0,1]^{|X|}$. In order to define uncertainty regions, it is natural then to consider hyperrectangles $\mathcal{I}\subseteq[0,1]^{|X|}$, i.e., subsets of the form =[$\mathcal{I}={}_{x\in X}[a_{x},b_{x}]$. Intuitively, these regions correspond to independently choosing a confidence interval $[a_{x},b_{x}]$ for the unknown probability $p_{x}$, for each feature $x\in X$.
@@
We can then reason about the importance of features via the analysis of SHAP polynomials. Here we concentrate on the fundamental problems of finding maxima and minima of these polynomials over the uncertainty region. This allows us to determine the SHAP interval of a feature, i.e., the set of possible SHAP scores of the feature over the uncertainty region. SHAP intervals may provide useful insights. For instance, obtaining smaller SHAP intervals for a feature suggests its SHAP score is more robust against uncertainty. On the other hand, they can be used to assess the relative importance of features (see Section 5 for more details).
@@
Problem: REGION-MAX-SHAP Input: A classifier $M$, an entity $e$, a feature $x$, a hyperrectangle $\mathcal{I}$ and a rational number $q$. Output: Is there a point $\textbf{p}\in\mathcal{I}$ such that $Shap_{M,e,x}(\textbf{p})\geq q$?.
The problem REGION-MIN-SHAP is defined analogously by requiring $Shap_{M,e,x}(\textbf{p})\leq q$. We also consider some related problems:
@@ 4.1 Preliminaries on multilinear polynomials
A vertex of a hyperrectangle =×$\mathcal{I}=\times_{i=1}^{n}[a_{i},b_{i}]$ is a point $\textbf{p}\in\mathcal{I}$ such that $\textbf{p}_{i}\in\{a_{i},b_{i}\}$ for each $1\leq i\leq n$. The following is a well-known fact about multilinear polynomials (see e.g.,  [14]). For completeness we provide a simple self-contained proof in Appendix A.1.
@@
Proposition 2 yields two algorithmic consequences. On the one hand, it induces an algorithm to find the maximum of $f$ over $\mathcal{I}$ in time $2^{n}poly(|f|)$: simply evaluate the polynomial on all the vertices, and keep the maximum [^5] We assume $f$ is given by listing its non-zero coefficients and their corresponding monomials.. On the other hand, it shows that this problem is certifiable: to decide whether $f$ can reach a value as big as $q$, we just need to guess the corresponding vertex and evaluate it. We show that, within the usual complexity theoretical assumptions, there is no polynomial algorithm to find this maximum (see Appendix A.1):
@@
Given a multilinear polynomial $f$, a hyperrectangle =[$\mathcal{I}={}_{i=1}^{n}[a_{i},b_{i}]$, and a rational $q$, the problem of deciding whether there is an $\textbf{x}\in\mathcal{I}$ such that $f(\textbf{x})\geq q$ is NP-complete.
@@ 4.2 Complexity of REGION-MAX-SHAP
It is well-known that computing SHAP scores is hard for general classifiers. For instance, the problem is already #P-hard when considering Boolean circuits [3] or logistic regression models [27]. For linear perceptrons, model counting is intractable [4] and by the results in [3], it follows that SHAP computation for perceptrons is also intractable.  On the other hand, computing maxima of SHAP polynomials for a certain class of classifiers is as hard as computing SHAP scores for that class of classifiers: if we consider the hyperrectangle consisting of a single point =[$\mathcal{I}={}_{i=1}^{n}[p_{i},p_{i}]$, the maximum of the SHAP polynomial coincides with the SHAP score for the product distribution $(p_{i})_{1\leq i\leq n}$. Therefore, we focus on family of classifiers where the SHAP score can be computed in polynomial time. A prominent example is the class of decomposable and deterministic Boolean circuits, whose tractable SHAP score computation has been shown recently [3]. This class contains as a special case the well-known class of decision trees. For a formal definition see  [10, 11]. As a consequence of Proposition 2, we obtain the following complexity upper bound:
@@
We reduce from the well-known NP-complete problem VERTEX-COVER: given a graph $G=(V,E)$ and $k\geq 1$, decide whether there is a vertex cover [^6] Recall a vertex cover of $G=(V,E)$ is a subset of the nodes $C\subseteq V$ such that for each edge $\{u,v\}\in E$, either $u\in C$ or $v\in C$. of size at most $k$.
@@
where $\ell$ is the size of the subset $C$ and the term $T_{n,\ell}$ grows as $\ell$ grows. The term $I_{u,v}$ is positive and works as a penalization factor which is “activated” if $p_{u}=p_{v}=1$, i.e., when the edge $uv$ is uncovered. Furthermore, by adding an extra feature $w$ to the model (and consequently, another probability $p_{w}$) we can make this penalization factor arbitrarily big in relation to the size factor $T_{n,\ell}$.
We choose the bound $q$ to be $-T_{n,k}$. If $C$ is a vertex cover of size $\ell$, then each term $p_{u}p_{v}I_{u,v}$ equals 0 and hence $Shap_{M,e,x}(\textbf{p}^{C})=-T_{n,\ell}$. On the other hand, if $C$ is not a vertex cover, then some $p_{u}p_{v}I_{u,v}$ is non-zero, and by defining an adequate interval for $p_{w}$ we can ensure that $-p_{v}p_{v}I_{u,v}<q$. Hence, the only way to obtain $Shap_{M,e_{0},x_{0}}(\textbf{p}^{C})\geq q$ is to pick $C$ to be, in the first place, a vertex cover, and secondly, one of size $\ell\leq k$. ∎